OpenStudy (tainted):

Please help me with these ( show work ) √(6x+5)=√53 and √(x+2+4)=x

OpenStudy (anonymous):

square both sides and solve \[6x+5=53\]

OpenStudy (anonymous):

\[(\sqrt{6x+5})^{2}= (\sqrt{53})^{2}\] By squaring both sides you get rid of the square roots: End up with: 6x+5=53. Then solve for x